OpenStudy (anonymous):

Solve the differential equation (dy/dx)=ycos(x)/(1+y^2) with the initial condition y(0) = 1.

OpenStudy (anonymous):

\[\frac{ dy }{dx }=\frac{ y \cos(x) }{ 1+y ^{2}}\]

OpenStudy (anonymous):

First we separate the values but idk how to do that in this case.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

if we separate the varibles, we get: \[dy\frac{{1 + {y^2}}}{y} = \cos x\;dx\]

OpenStudy (michele_laino):

variables*

OpenStudy (anonymous):

@Michele_Laino then we intergrate can get this \[\int\limits \frac{ 1+y ^{2} }{ y } dy= \int\limits \cos(x) dx\] \[\ln(y)+\frac{ y ^{2} }{ 2}+C=\sin(x)+C\] \[\ln(y) + \frac{ y ^{2} }{ 2 }=\sin(x)+C\]

OpenStudy (anonymous):

@Michele_Laino what then?

OpenStudy (michele_laino):

we have to set y=1 and x=0, so we can find the value of the constant C

OpenStudy (anonymous):

okay so plug in the x and y values or should we simplify the rest of the equation first. can if we have to simplify what would i do next?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

it is the same, we can rewrite your solution as below: \[\Large y{e^{{y^2}/2}} = C{e^{\sin x}}\]

OpenStudy (michele_laino):

substantially it is the same as yours

OpenStudy (michele_laino):

if I set x=0 and y=1, I get this: \[\Large C = \sqrt e \]

OpenStudy (anonymous):

Yes! I got the same answer so now what do we do? @Michele_Laino

OpenStudy (anonymous):

plug sqrt(e) back into the equation

OpenStudy (anonymous):

?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

after that substitution, we get: \[\Large y{e^{{y^2}/2}} = \sqrt e {e^{\sin x}}\]

OpenStudy (anonymous):

\[ye ^{\frac{ y ^{2} }{ 2 }}=(\sqrt{e})e ^{\sin(x)}\]

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

@Michele_Laino is that it then?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

thanks!!

OpenStudy (michele_laino):

thanks! :)